OpenStudy (anonymous):

how do you find the inverse of y=x^2 +2

OpenStudy (dumbcow):

get x by itself y-2 = x^2 sqrt(y-2) = x

OpenStudy (anonymous):

i thought you had to switch x and y. so you would have x=y^2 +2 then you subtract 2 so its x-2=y^2

OpenStudy (dumbcow):

either way works you will still get f^-1 = sqrt(x-2)

OpenStudy (anonymous):

so the answer is ?

OpenStudy (anonymous):

y= sqrt of x-2?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

thank you. can you help me some more.?

OpenStudy (dumbcow):

what do you have

OpenStudy (anonymous):

the direcxtions say find ^-1 and the doman and range of f (-1) determine where f^-1 is a function. and the problem is f(x) sqrt 3x

OpenStudy (dumbcow):

ok flip x and y x = sqrt(3y) solve for y

OpenStudy (anonymous):

so the anser is x^2 = y

OpenStudy (dumbcow):

dont forget about the 3 x^2 = 3y (x^2/3) = y now what is domain and range

OpenStudy (anonymous):

domain is all real. and range idk

OpenStudy (dumbcow):

good range is all possible y-values since x is squared, all the negative x's we put in become positive right? so our y value is always positive

OpenStudy (anonymous):

ohh. so its y is greater than 0 ?

OpenStudy (dumbcow):

correct, well equal to y>=0

OpenStudy (anonymous):

oh em gee thank you. :]

OpenStudy (dumbcow):

your welcome